OpenStudy (anonymous):

Ok, 1 more! 7th grade math A political party estimates that it's candidate will receive between 28% and 33% of the votes during an election. If 639 people vote, about how many votes can the candidate expect to receive?

OpenStudy (anonymous):

@Lovelarap

OpenStudy (anonymous):

@ChantySquirrel1129**

OpenStudy (anonymous):

Convert you percent in to numbers so now you have 0.28 and 0.33. Now multiply it by 639 and you should get an answer. : )

OpenStudy (anonymous):

Both of the percents? how will I know which one is correct?

OpenStudy (anonymous):

cause it's saying "between" so the numbers you get will be "between" the two percents.

OpenStudy (compassionate):

\[(628 \times 0.22 \space \space \space \space, \space \space \space \space 628 \times 0.33)\]

OpenStudy (anonymous):

wait 22 now?

OpenStudy (compassionate):

Oopsies. Haha. I meant 0.28 and 0.33.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

ok, I got 1936.36 for 0.33 and 0.28 2282.14

OpenStudy (anonymous):

so what now?

OpenStudy (anonymous):

those are your answers, if i'm not wrong

OpenStudy (anonymous):

it says about how many votes can the candidate see

OpenStudy (compassionate):

Now just find the average. Add the numbers together and divide by the number of numbers. (The sum of values divided by the number of values). Arrange the numbers in order, find the middle number. (The middle value when the values are ranked).

OpenStudy (anonymous):

expect to receive* jeez i'm tired

OpenStudy (compassionate):

And don't use decimals. It doesn't make since to have .18th of a vote.

OpenStudy (anonymous):

wait so I add 936 and 2282?

OpenStudy (compassionate):

\[average = \frac{ Votes }{ Number \space of \space votes }\]

OpenStudy (anonymous):

1936*

OpenStudy (compassionate):

Yes. Add them, then divide it by the number of votes (since they received .28 and .33, the number of votes are 2.)

OpenStudy (anonymous):

I got 6.6

OpenStudy (anonymous):

incorrect?

OpenStudy (anonymous):

or 6

OpenStudy (anonymous):

wait no 7

OpenStudy (anonymous):

eh?

OpenStudy (compassionate):

No. Add the numbers together then divide it by 2

OpenStudy (compassionate):

Do you understand what average is?

OpenStudy (anonymous):

thats what I did

OpenStudy (anonymous):

pls help me get the answer quick it's way pass my bedtime

OpenStudy (anonymous):

already almost 1 in the morning

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

\[\frac{ 936 + 2282 }{ 2 }\]

OpenStudy (anonymous):

oh ty

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

I got 2109

OpenStudy (anonymous):

hm? i got 1609

OpenStudy (compassionate):

2109 is correct. Goodnight @LoveIt

OpenStudy (anonymous):

oh, ok gn

OpenStudy (compassionate):

I don't feel you learned anything from this, however.

OpenStudy (compassionate):

I feel you just want an answer and aren't thinking about it. But, whatever. You do you, boo.

OpenStudy (anonymous):

I'll review this tomorrow, it's just that i'm really tired sorry

OpenStudy (anonymous):

k gn bye